# In-vitro measurements coupled with in-silico simulations for stochastic calibration and uncertainty quantification of the mechanical response of biological materials

### v1

@@ 2.4 Bayesian optimization and sensitivity analysis
To determine specimen-specific biomechanical parameters, and use the minimum and maximum bounds of these parameters for uncertainty quantification scheme, Bayesian optimization (BO) is implemented using scikit-optimize library. BO is an iterative framework focused on optimizing an expensive and unknown objective function $f(x)$, where $x\in\mathbb{R}^{d}$ is a vector of $d$ parameters. The goal is to solve
@@
where $\mathcal{A}$ is the search space. The gradient information is unavailable for $f(x)$ and is considered a “black box” function. The core of BO is based on constructing a probabilistic surrogate model, a Gaussian process (GP) to approximate $f(x)$. It models the unknown objective function $f(x)$ by providing a posterior distribution given the observed data. With its simplest form, the resulting prior of GP is written as:
@@
where $\mu(x)$ is the mean function, showing the prior belief about the objective function and $k(x,x^{\prime})$ is the covariance (or kernel) function. Once building the surrogate model, Bayesian optimization selects the next point $x_{\text{next}}$ to evaluate by maximizing an acquisition function. A common acquisition function used in scikit-optimize is Expected Improvement (EI) and defined as:
@@
where $f^{*}$ is the current best observed value of $f(x)$. The $f(x)$ function at any $x$ point is normally distributed with mean $\mu(x)$ and variance $\sigma^{2}(x)$, and the improvement is modeled as:
@@
where $\Phi(\cdot)$ is the cumulative distribution function (CDF) and $\phi(\cdot)$ is the probability density function (PDF) of the standard normal distribution, respectively. This acquisition function seeks regions where the model predicts the largest improvement over the current best value. To find the next sampling point $x_{\text{next}}$, the acquisition function $\text{EI}(x)$ is maximized. Once a new sample $x_{\text{next}}$ is selected and the objective function $f(x_{\text{next}})$ is evaluated, the surrogate model (the Gaussian Process) is updated. The process continues until a stopping criterion is met. Further details of BO are documented in literature [frazier2018tutorialbayesianoptimization]. We have made a couple of trials, and observed that a stop criteria with 100 fixed number iterations is reasonable for tuning the biomechanical parameters. We note that a higher number of iterations increased the computational cost. The objective function to be minimized is defined as the summation of the root mean square error (RMSE) between experimental (in-vitro) and numerical (in-silico) total reaction force values for the three loading scenarios. It is given by:
[…]
@@
where $R_{ij}$ and $\mathcal{S}{(u_{ij},x})$ are the experimental and numerical reaction forces, respectively, for the parameter set $x$ at displacement $u_{ij}$, and data point i. The term $n_{j}$ denotes the number of data points for test loading j, where j = 1 corresponds to the one-step relaxation, j = 2 to the two-step relaxation, and j = 3 to the incremental cyclic loading scenario. The factor of 2 in the equation arises from the symmetric boundary conditions modeled in the finite element (FE) simulation. Simply, the $(R_{ij},u_{ij})$ are in-vitro observation pairs, and $\mathcal{S}(u_{ij},x)$ is the unknown force function that extracts the corresponding data from the FE simulator. The BO tries to find the model parameters $x$ that minimize the function $f(x)$ by repeatedly evaluating the function $\text{EI}(x)$, which calls the simulator (FEBio) to solve the black-box problem. The parameter search ranges used during BO are given in Table  1, and the initial search ranges for these parameters are predicted as follows. First, the viscoelastic characteristics $g_{i}$ and $\tau_{i}$ are estimated by fitting on some trial one-step relaxation measurements using the Levenberg-Marquardt method. Then, the bounds of the estimated parameters are extended and used as search ranges for BO. Similar search ranges are considered for the fiber and matrix components. For the approximate prediction of the modulus, we initially evaluated it by analyzing the linear region of a typical incremental cyclic test loading measurement and found that the modulus of elasticity for the composite tendon (matrix + fiber) is around 350 MP. Although, to the best of our knowledge, no study has experimentally quantified the elastic modulus of the fiber and elastic component in the human Achilles tendon separately, it has been stated in some rat studies that the fiber-to-matrix ratio is between 5 and 10. [dutov_measurement_2016, khayyeri_fibre-reinforced_2015, lewis_modeling_1997]. We considered these uncertain ranges during the determination of the search ranges.
@@
Sensitivity and feature importance analyses are performed to investigate the influence of each parameter on the FRVE model for each test loading scenario. Feature importance is determined using the random forest algorithm implemented in scikit-learn, which is based on impurity reduction. The importance of a feature is computed as the total reduction of the criterion observed by the feature, known as the Gini importance [breiman_classificatin]. The importance of a feature $I(x_{i})$ is given by:
@@
where $n_{t}$ is the number of trees in the forest, $S_{t}(x_{i})$ is the set of all splits for the feature $x_{i}$ in the tree $t$, and $\Delta I_{s}$ is the decrease in the splitting criteria at split $s$. The final importance values are then normalized to 100 over all features. More details on the random forest algorithm are documented in the literature [breiman_classificatin].
As an addendum to the feature importance analysis, we implemented Sobol sensitivity analysis to investigate the sensitivity of the input parameters on the model output using salib [herman_salib_2017]. It is a variance-based approach to quantify the parameter’s sensitivity as well as their contribution to the model. Basically, it decomposes the total variance of the output $Y$ into contributions that are attributable to input parameters and their interactions. Let us assume that $Y=f(X_{1},X_{2},\dots,X_{d})$ represents the model output as a function of $d$ input parameters $X_{1},X_{2},\dots,X_{d}$. The total variance of $Y$ is given by:
[…]
@@
where $V_{i}=\mathrm{Var}\left(E[Y\mid X_{i}]\right)$ represents the main effect of input $X_{i}$, and $V_{ij}=\mathrm{Var}\left(E[Y\mid X_{i},X_{j}]\right)-V_{i}-V_{j}$ shows the interaction effect between $X_{i}$ and $X_{j}$. Higher-order terms denote interactions between more input parameters. The Sobol indices are dimensionless quantities that measure the relative contribution of each term to the total variance. The first order sensitivity index is given by:
$S_{i}=\frac{V_{i}}{V}$ (13)
The 1st order index shows the proportion of variance that is attributed to $X_{i}$. Total-order sensitivity index is given by:
@@
where $V_{\sim i}$ is the variance of the output excluding $X_{i}$. The total-order index accounts for both the main effect and all interaction effects involving $X_{i}$. The interaction terms (e.g., $S_{ij}$) can be calculated similarly by dividing the corresponding variance term by the total variance. However, the sensitivity due to interaction is not covered in this study. We implemented Saltelli sampling to create the parameter space of 20,000 points for each of the 16 input parameters of the FRVE model. The outputs are the results computed from in-silico simulations. The sampling ranges used for BO and ABC are also used during Saltelli sampling (Table  1). We used the minimization function values given in Eq. 9 for j = 1, 2, 3, as the results of the Sobol sensitivity analysis. More details on the Sobol sensitivity analysis are reported in the previous study [SALTELLI2010259].
@@ 2.5 Approximate Bayesian Computation
Although the standard Bayesian framework allows for updating prior beliefs about parameters based on the likelihood of the data given those parameters, evaluating the likelihood $P(\mathbf{y}_{\text{obs}}|\boldsymbol{\theta})$ is often infeasible, specifically for complex models, but the model can generate synthetic data $\mathbf{y}_{\text{sim}}$ given parameters $\boldsymbol{{\theta}}$. We implemented likelihood-free approximate Bayesian computation (ABC) for stochastic calibration and uncertainty quantification of posterior parameters using PyABC [schalte_pyabc_2022]. The framework is based on an executable “black-box” forward process model that simulates data using given model parameters [Price02012018, wilkonson_2013]. In ABC, the computation of the likelihood is replaced by the comparison between observed (in-vitro measurements) $\mathbf{y}_{\text{obs}}$ and simulated data (in-silico simulations) $\mathbf{y}_{\text{sim}}$. Given the prior distribution of $P(\boldsymbol{\theta})$ of parameters $\boldsymbol{\theta}$, the target is to approximate the posterior distribution $P(\boldsymbol{\theta}|\mathbf{y}_{\text{obs}})$ given observed data $\mathbf{y}_{\text{obs}}$, a model $f(\mathbf{y}_{\text{sim}}|\boldsymbol{\theta})$, and a prior distribution $P(\boldsymbol{\theta})$. Typically, it is based on generating simulated data $\mathbf{y}_{\text{sim}}$ for parameters $\boldsymbol{\theta}$ drawn from the prior distribution and retaining parameters for which $\mathbf{y}_{\text{sim}}$ is sufficiently close to $\mathbf{y}_{\text{obs}}$. The posterior is approximated as:
@@
where $d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})$ is a distance metric, $\epsilon$ is a tolerance threshold, and $\mathbb{I}(\cdot)$ is an indicator function that equals 1 when the condition is satisfied $d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})\leq\epsilon$ and 0 otherwise $d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})\geq\epsilon$. The output of the algorithm is a sample set of parameters from the distribution $P(\theta|d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})\leq\epsilon)$. Direct ABC implementation can be inefficient, particularly for small ϵ, as it reduces the number of accepted parameters. To overcome this, pyABC performs a Sequential Monte Carlo (SMC) technique that progressively refines the posterior estimation. It facilitates robust and efficient inference for a broad spectrum of applications. Further detail of ABC model is documented in previous studies [toni_simulation-based_2010].
In this study, the prior distributions for the parameters are defined based on a uniform distribution (Table  1). The distance between observed and simulated data is computed using the Euclidean distance (L2) metric. To refine the parameter estimates, the $\epsilon$ tolerance was adaptively adjusted in the ABC SMC process. The initial tolerance threshold of $\epsilon$ is set as 0.3. We note that we selected these options by running some trials and focusing on where the simulated data $\mathbf{y}_{\text{sim}}$ most closely matched the observed data $\mathbf{y}_{\text{obs}}$
@@ 2.6 Implementation procedures
The implementation begins with specimen-specific parameter calibration to determine the approximate values of the FRVE parameter set using an optimization technique. We employed Bayesian Optimization (BO), and the procedures and stages of its implementation are illustrated in Figure  2, where the numbers indicate data processing steps. Initially, in-vitro test samples are harvested and re-sectioned (1). Then in vitro test loading protocols are applied and their measurements are acquired, categorized and saved as a serialized dataset file (2). As a preliminary study, we chose a simple uniaxial tensile test for the three test loading scenarios. For each test, the reaction force $R(t)$ is obtained as a function of displacement $u(t)$. This process is applied to all samples. Next, the search space is defined, and the objective function is formulated (3). In-silico simulations using the finite element method are performed to obtain time vs. reaction force values (4). The objective function is computed as the difference between in-silico and in-vitro reaction force observations (5). Once the objective function is minimized in the current step, the next target is to optimize the acquisition function, balancing exploration and exploitation (6-7). The model is then updated by incorporating a new observation and retraining the surrogate model. Finally, steps 3-7 are iterated until the stopping criteria are met (8). To reduce computational time, we set the stopping criterion at 100 iterations, as this was sufficient for calibrating specimen-specific parameters. The implementation of in-vitro measurements, coupled with in-silico simulation and ABC, resembles the BO stages at the pre-processing steps. However, the main difference is that instead of evaluating specimen-specific observations, ABC includes all specimens’ measurements, allowing for the stochastic calibration of parameters, including uncertainty. Further details are as follows (Figure 3).
@@
    The importance and sensitivity of each parameter are determined using the Random Forest and Sobol methods. Once the minimum and maximum bounds obtained from BO across all samples are determined, these wide bounds are used as priors in ABC. Initially, parameters $\boldsymbol{\theta}$ are drawn from the prior uniform distribution as: $\boldsymbol{\theta}\sim P(\boldsymbol{\theta})$.
@@
    The next step is to perform forward simulations to drive ABC. We implemented FEBio as the simulator and created a template that includes the FRVE material model and boundary conditions that mimic the experiment. FEBio input files (.feb) are xml type files, and we used the jinja package to create the FEBio input template file and serialize the data between ABC and the simulator. Once ABC determines the input parameter set $\boldsymbol{\theta}$, it is passed directly to this simulator template, and the FEBio simulator generates simulated data $f(\mathbf{y}|\boldsymbol{\theta})$ using $\boldsymbol{\theta}$ as: $\mathbf{y}_{\text{sim}}\sim f(\mathbf{y}|\boldsymbol{\theta})$.
@@
    The parameter $\boldsymbol{\theta}$ is retained if the distance satisfies: $d\leq\epsilon$; otherwise, return to step 4.

### v2

@@ 2.4 Bayesian optimization and sensitivity analysis
To determine specimen-specific biomechanical parameters, Bayesian optimization (BO) is implemented. BO was selected primarily due to the high computational cost of finite element simulations involved in evaluating the objective function $f(x)$. Each evaluation requires solving a full-scale FE model, making traditional methods such as grid search or evolutionary algorithms computationally impractical. Beyond efficient parameter calibration, BO also serves to inform the parameter set values used as priors in the ABC-based uncertainty quantification framework. It is an iterative framework focused on optimizing an expensive and unknown objective function $f(x)$, where $x\in\mathbb{R}^{d}$ is a vector of $d$ parameters. The goal is to solve
@@
where $\mathcal{A}$ is the search space. Gradient information is not available for $f(x)$ and is considered a “black box” function. The core of BO is based on the construction of a probabilistic surrogate model, a Gaussian process (GP), to approximate $f(x)$. It models the unknown objective function $f(x)$ by providing a posterior distribution given the observed data. With its simplest form, the resulting prior of GP is written as:
@@
where $\mu(x)$ is the mean function, showing the prior belief in the objective function and $k(x,x^{\prime})$ is the covariance function (or kernel). In this study, the surrogate model was built using a Gaussian Process (GP) with a Matérn kernel, as provided by the scikit-optimize library. An anisotropic kernel was used, allowing a separate length-scale to be estimated for each parameter dimension. The optimization process was initialized using a random sampling strategy. Once building the surrogate model, Bayesian optimization selects the next point $x_{\text{next}}$ to evaluate by maximizing an acquisition function. A common acquisition function used in scikit-optimize is Expected Improvement (EI) and is defined as:
@@
where $f^{*}$ is the best observed current value of $f(x)$. The function $f(x)$ at any $x$ point is normally distributed with mean $\mu(x)$ and variance $\sigma^{2}(x)$, and the improvement is modeled as:
@@
where $\Phi(\cdot)$ is the cumulative distribution function (CDF) and $\phi(\cdot)$ is the probability density function (PDF) of the standard normal distribution, respectively. This acquisition function seeks regions where the model predicts the largest improvement over the current best value. To find the next sampling point $x_{\text{next}}$, the acquisition function $\text{EI}(x)$ is maximized. Once a new sample $x_{\text{next}}$ is selected and the objective function $f(x_{\text{next}})$ is evaluated, the surrogate model (the Gaussian Process) is updated. The process continues until a stopping criterion is met. Further details of BO are documented in literature [38]. In preliminary tests, threshold-based stopping led to extended runtimes without significant gains in accuracy. Therefore, we adopted a fixed-iteration approach of 100 iterations for practical and computational efficiency. The objective function to be minimized is defined as the summation of the root mean square error (RMSE) between experimental (in-vitro) and numerical (in-silico) total reaction force values for the three loading scenarios. It is given by:
$\text{$f(x)$}=\sum_{j=1}^{3}\sqrt{\frac{1}{n_{j}}\sum_{i=1}^{n_{j}}\left[R_{ij}-2\mathcal{S}{(u_{ij},x})\right]^{2}}$ (9)
@@
where $R_{ij}$ and $\mathcal{S}{(u_{ij},x})$ are the experimental and numerical reaction forces, respectively, for the parameter set $x$ at displacement $u_{ij}$, and data point i. The term $n_{j}$ denotes the number of data points for test loading j, where j = 1 corresponds to the one-step relaxation, j = 2 to the two-step relaxation, and j = 3 to the incremental cyclic loading scenario. The factor of 2 in the equation arises from the symmetric boundary conditions modeled in the finite element (FE) simulation. Simply, the $(R_{ij},u_{ij})$ are in-vitro observation pairs, and $\mathcal{S}(u_{ij},x)$ is the unknown force function that extracts the corresponding data from the FE simulator. The BO tries to find the model parameters $x$ that minimize the function $f(x)$ by repeatedly evaluating the function $\text{EI}(x)$, which calls the simulator (FEBio) to solve the black-box problem. The parameter search ranges used during BO are given in Table  1, and the initial search ranges for these parameters are predicted as follows. First, the viscoelastic characteristics $g_{i}$ and $\tau_{i}$ are estimated by fitting on some trial one-step relaxation measurements using the Levenberg-Marquardt method. Then, the bounds of the estimated parameters are extended and used as search ranges for BO. Similar search ranges are considered for the fiber and matrix components. For the approximate prediction of the modulus, we initially evaluated it by analyzing the linear region of a typical incremental cyclic test loading measurement and found that the modulus of elasticity for the composite tendon (matrix + fiber) is around 350 MP. Although, to the best of our knowledge, no study has experimentally quantified the elastic modulus of the fiber and elastic component in the human Achilles tendon separately, it has been stated in some rat studies that the fiber-to-matrix ratio is between 5 and 10 [39, 40, 41].
@@
Sensitivity and feature importance analyzes are performed to investigate the influence of each parameter on the FRVE model for each test loading scenario. Feature importance is determined using the random forest algorithm implemented in scikit-learn, which is based on impurity reduction. The importance of a feature is computed as the total reduction of the criterion observed by the feature, known as the Gini importance [42]. The importance of a feature $I(x_{i})$ is given by:
@@
where $n_{t}$ is the number of trees in the forest, $S_{t}(x_{i})$ is the set of all splits for the feature $x_{i}$ in the tree $t$, and $\Delta I_{s}$ is the decrease in the splitting criteria at split $s$. The final importance values are then normalized to 100 over all features. More details on the random forest algorithm are documented in the literature [42].
As an addendum to the feature importance analysis, we implemented Sobol sensitivity analysis to investigate the sensitivity of the input parameters on the model output using salib [43]. It is a variance-based approach to quantify the parameter’s sensitivity as well as their contribution to the model. Basically, it decomposes the total variance of the output $Y$ into contributions that are attributable to input parameters and their interactions. Let us assume that $Y=f(X_{1},X_{2},\dots,X_{d})$ represents the model output as a function of $d$ input parameters $X_{1},X_{2},\dots,X_{d}$. The total variance of $Y$ is given by:
$V=\mathrm{Var}(Y)=\int\left(f(\mathbf{X})-E[Y]\right)^{2}p(\mathbf{X})\,d\mathbf{X}$ (11)
@@
where $V_{i}=\mathrm{Var}\left(E[Y\mid X_{i}]\right)$ represents the main effect of input $X_{i}$, and $V_{ij}=\mathrm{Var}\left(E[Y\mid X_{i},X_{j}]\right)-V_{i}-V_{j}$ shows the interaction effect between $X_{i}$ and $X_{j}$. Higher-order terms denote interactions between more input parameters. The Sobol indices are dimensionless quantities that measure the relative contribution of each term to the total variance. The first-order sensitivity index is given by:
$S_{i}=\frac{V_{i}}{V}.$ (13)
The first-order index shows the proportion of variance that is attributed to $X_{i}$. Total-order sensitivity index is given by:
@@
where $V_{\sim i}$ is the variance of the output excluding $X_{i}$. The total-order index accounts for both the main effect and all interaction effects involving $X_{i}$. The interaction terms (e.g., $S_{ij}$) can be calculated similarly by dividing the corresponding variance term by the total variance. However, the sensitivity due to interaction is not covered in this study. We implemented Saltelli sampling to create the parameter space of 20,000 points for each of the 16 input parameters of the FRVE model. In our Saltelli-based Sobol sensitivity analysis, we employed uniform distributions over the parameter ranges as defined in Table  1. The model output used for Sobol analysis was the total error function defined in Eq. 9 which aggregates discrepancies across three experimental loading scenarios. Additional details on the Sobol sensitivity analysis are reported in the previous studies [44, 45, 46].
@@ 2.5 Approximate Bayesian Computation
Although the standard Bayesian framework allows for updating prior beliefs about parameters based on the likelihood of the data given those parameters, evaluating the likelihood $P(\mathbf{y}_{\text{obs}}|\boldsymbol{\theta})$ is often infeasible, but the model can generate synthetic data $\mathbf{y}_{\text{sim}}$ given parameters $\boldsymbol{{\theta}}$. In general, Bayesian inference provides a principled framework for quantifying uncertainty in model parameters by combining prior knowledge with evidence from observed data through the likelihood function. However, for many nonlinear models, deriving an analytical or tractable likelihood function is difficult or impossible. Here, we implemented likelihood-free approximate Bayesian computation (ABC) for stochastic calibration and uncertainty quantification of posterior parameters using PyABC [31]. The framework is based on an executable “black-box” forward process model that simulates data using given model parameters [47, 48]. In ABC, the computation of the likelihood is replaced by the comparison between observed (in-vitro measurements) $\mathbf{y}_{\text{obs}}$ and simulated data (in-silico simulations) $\mathbf{y}_{\text{sim}}$. Given the prior distribution of $P(\boldsymbol{\theta})$ of parameters $\boldsymbol{\theta}$, the target is to approximate the posterior distribution $P(\boldsymbol{\theta}|\mathbf{y}_{\text{obs}})$ given observed data $\mathbf{y}_{\text{obs}}$, a model $f(\mathbf{y}_{\text{sim}}|\boldsymbol{\theta})$, and a prior distribution $P(\boldsymbol{\theta})$. Typically, it is based on generating simulated data $\mathbf{y}_{\text{sim}}$ for parameters $\boldsymbol{\theta}$ drawn from the prior distribution and retaining parameters for which $\mathbf{y}_{\text{sim}}$ is sufficiently close to $\mathbf{y}_{\text{obs}}$. The posterior is approximated as:
@@
where $d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})$ is a distance metric, $\epsilon$ is a tolerance threshold, and $\mathbb{I}(\cdot)$ is an indicator function that equals 1 when the condition is satisfied $d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})\leq\epsilon$ and 0 otherwise $d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})\geq\epsilon$. The output of the algorithm is a sample set of parameters from the distribution $P(\theta|d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})\leq\epsilon)$. Direct implementation of ABC can be inefficient, particularly for small $\varepsilon$, as it reduces the number of accepted parameters. To overcome this, pyABC performs a Sequential Monte Carlo (SMC) technique that progressively refines the posterior estimation. It facilitates robust and efficient inference for a broad spectrum of applications. Further details of the ABC model are documented in previous studies [30].
In this study, the prior distributions for the model parameters are defined based on the results of specimen-specific calibrations using Bayesian optimization (Table 2). Specifically, the mean and standard deviation of each parameter, computed across nine specimens for each of the three sites, are used to construct independent Gaussian priors. To account for a potential underestimation of biological variability due to the point-estimate nature of Bayesian optimization, and to ensure coverage of physiologically plausible ranges, the standard deviations are conservatively inflated by a factor of two. The distance between observed and simulated data is computed using the Euclidean distance (L2) metric. To refine parameter estimates, the tolerance $\epsilon$ is adaptively adjusted in the ABC SMC process. The initial tolerance threshold of $\epsilon$ is set as 0.3. We note that we selected these options by running some trials and focusing on where the simulated data $\mathbf{y}_{\text{sim}}$ most closely matched the observed data $\mathbf{y}_{\text{obs}}$.
@@ 2.6 Implementation procedures
The implementation begins with the calibration of the specimen-specific parameters to determine the approximate values of the set of FRVE parameters using an optimization technique. We employed Bayesian Optimization (BO), and the procedures and stages of its implementation are illustrated in Figure  2, where the numbers indicate data processing steps. Initially, test samples are harvested and re-sectioned (1). Then in-vitro test loading protocols are applied and their measurements are acquired, categorized and saved as a serialized dataset file (2). As a preliminary study, we chose a simple uniaxial tensile test for the three test loading scenarios. For each test, the reaction force $R(t)$ is obtained as a function of displacement $u(t)$. This process is applied to all samples. Next, the search space is defined, and the objective function is formulated (3). In-silico simulations are performed using the finite element method to obtain time vs. reaction force values (4). The objective function is computed as the difference between in-silico and in-vitro reaction force observations (5). Once the objective function is minimized in the current step, the next target is to optimize the acquisition function, balancing exploration and exploitation (6-7). The model is then updated by incorporating a new observation and retraining the surrogate model. Finally, steps 3-7 are iterated until the stopping criteria are met (8). The implementation of in-vitro measurements, coupled with in-silico simulation and ABC, resembles the BO stages at the pre-processing steps. However, the main difference is that instead of evaluating specimen-specific observations, ABC includes all specimens’ measurements, allowing for the stochastic calibration of parameters, including uncertainty. The implementation procedure and sequence of steps illustrated in Figure 3 are as follows:
@@
    The importance and sensitivity of each parameter are assessed using Random Forest and Sobol methods. The model parameters obtained from Bayesian optimization (BO) across all samples are then used to inform the priors in the ABC framework.
@@
    The next step is to perform forward simulations to drive ABC. We implemented FEBio as the simulator and created a template that includes the FRVE material model and boundary conditions that mimic the experiment. FEBio input files (.feb) are xml type files, and we used the jinja package to create the FEBio input template file and serialize the data between ABC and the simulator. Once ABC determines the input parameter set $\boldsymbol{\theta}$ that drawn from the distribution, it is passed directly to this simulator template, and the FEBio simulator generates simulated data $f(\mathbf{y}|\boldsymbol{\theta})$ using $\boldsymbol{\theta}$ as: $\mathbf{y}_{\text{sim}}\sim f(\mathbf{y}|\boldsymbol{\theta})$.
@@
    The parameter set $\boldsymbol{\theta}$ is retained if the distance satisfies: $d\leq\epsilon$; otherwise, return to step 4.